# Open transmission channels in multimode fiber cavities with random mode mixing

### v1

@@ Abstract
The transport of light in disordered media is governed by open transmission channels, which enable nearly complete transmission of the incident power, despite low average transmission. Extensively studied in diffusive media and chaotic cavities, open channels exhibit unique properties such as universal spatial structure and extended dwell times. However, their experimental study is challenging due to the large number of modes required for control and measurement. We propose a multimode fiber cavity (MMFC) as a platform to explore open channels. Leveraging mode confinement and finite angular spread, MMFCs enable full control over all channels. This allowed us to achieve an 18-fold power enhancement by selectively exciting an open channel with a transmission rate of $0.90\pm 0.04$. By analyzing 100 transmission matrices of MMFC realizations, we observed a bimodal transmission eigenvalue distribution, indicating high channel control and low losses. The scalability of MMFCs, combined with long dwell times and potential for nonlinear phenomena, offers new opportunities for studying complex wave transport.
@@
Transmission through multimode optical fibers shares similarities with scattering samples, as both exhibit random mode coupling [18, 19, 20]. In standard optical fibers, open channels have little significance since backscattering is negligible. However, by introducing reflective coatings to create a multimode fiber cavity (MMFC), most incident light is backreflected and the concept of open channels becomes applicable. Without mode mixing, an open channel of an MMFC simply corresponds to a guided mode whose propagation constant matches the Fabry–Pérot resonance condition. However, in the presence of strong mode mixing, the existence and properties of open channels ̵ֿare more intricate.
@@
Open channels are found by measuring the transmission matrix of the sample $T$, which relates the input and output fields by $\vec{\psi}_{out}=T\vec{\psi}_{in}$ [23]. The singular value decomposition of $T$, or equivalently, the eigenvectors of the Hermitian matrix $T^{\dagger}T$, define the transmission eigenchannels of the samples $T^{\dagger}T\vec{v}_{n}=\tau_{n}\vec{v}_{n}$, where the eigenvalues $\tau_{n}$ correspond to the transmission rates. To selectively excite the most open channel, the wavefront of the input field is tailored to match the wavefront corresponding to the eigenchannel $\vec{v}_{1}$ with the highest transmission rate $\tau_{1}$. The two key ingredients for realizing open channels are, therefore, coherent detection of the light transmitted through the fiber and precise control over the amplitude and phase of the incident wavefront.
@@
To experimentally study the transmission eigenchannels of an MMFC, we used a one-meter long step-index fiber with a core diameter of $25\,{\mathrm{\mu m}}$ and a numerical aperture of NA = 0.1, coated with reflective coatings with reflectivity $\rho=0.88$. We placed the MMFC in one arm of a Mach-Zehnder interferometer (Fig. 1a and Fig. S1). We coupled to the interferometer a tunable laser ($\lambda=632\,\mathrm{nm}$) with a long coherence length ($>$ 100 m) which allowed interference of the multiple round trips of the light in the cavity. The amplitude and phase of the incident field were tailored using a phase-only spatial light modulator (SLM) in combination with a spatial filter for amplitude modulation [24] (see Supplementary Material, Section I.1). At the output of the fiber, the transmitted field was interfered with a reference beam in an off-axis holography configuration, and the two output polarizations were imaged onto a camera. We normalized the reconstructed output field so that its total power matched the output power measured by a calibrated photodiode monitoring the transmitted light (see Supplementary Material Fig. S1 and Section I.2).
@@
The eigenvalues obtained by diagonalizing $T^{\dagger}T$ are presented in Fig. 2. We observed five channels with transmission rates higher than four times the average transmission, $\langle\tau\rangle=0.064\pm 0.001$. The presence of multiple channels with transmission rates exceeding $4\langle\tau\rangle$ demonstrates significant control over the incident wavefront. In contrast, partial control results in a transmission rate distribution that rapidly converges to the Marčenko-Pastur distribution, for which the maximal transmission rate for square matrices is $4\langle\tau\rangle$ [23, 25]. Additionally, we observed a consistent decrease in the eigenvalues up to a pronounced gap at the 68th eigenvalue. This suggests that the MMFC supports 68 modes, in agreement with the measured transmission matrix of the fiber without dielectric coatings, which exhibits 34 guided modes and two leaky modes per polarization (see Supplementary Material Section I.5).
Figure 2: Transmission rates $\tau_{n}$ obtained by diagonalizing $T^{\dagger}T$ for an experimentally measured transmission matrix $T$. Only the 70 most significant transmission rates are presented. The gap between channels 68 and 69 indicates that the MMFC supports 68 guided modes. The dashed horizontal line marks the average transmission $\langle\tau\rangle=0.064\pm 0.001$. The uncertainties are in the order of $1\%$, smaller than the marker size (see Supplementary Material Section I.4)
@@
Figure 4: Transmission matrix and mode mixing in a multimode fiber cavity (MMFC). (a) Schematic of the MMFC: Each round trip is described by $\hat{r}_{1}T_{0}^{\mathsf{T}}\hat{r}_{2}T_{0}$, where $T_{0}$ and $T_{0}^{\mathsf{T}}$ are the transmission matrix of an uncoated fiber and its transpose, respectively, and $\hat{r}_{1},\hat{r}_{2}$ denote the matrices describing the reflection from the facets. (b) Typical measured transmission matrix $T_{0}$ of a one-meter long uncoated stepped-index fiber supporting $N=34$ modes per polarization. The matrix exhibits weak mode coupling, with 0.8 of the energy concentrated in the diagonal blocks of the degenerate modes. (c) Insets show numerically computed reflection matrices obtained by decomposing a linear phase tilt into the fiber modes. (d) Computed transmission matrix of the MMFC using Eq. (1), assuming facet angles of $\theta=5\times 10^{-3}$ rad.
@@
The slight deviation of the measured and model PDFs can be used to estimate the degree of control over the MMFC channels, as incomplete wavefront control or system losses typically reduce the number of open channels. We therefore considered two simplified single-parameter models. In the incomplete channel control (ICC) model, we assumed a lossless MMFC, while allowing for an incomplete measurement of its transmission matrix (see Supplementary Material Section I.7.2 and Ref.[25]). The best agreement between the ICC model and the experimental data was obtained by assuming full control over all incident channels and 95% control over the output channels (Fig. 2, orange curve). In the loss model, we assumed the MMFC is lossy due to leaky fiber modes but that its transmission matrix was perfectly measured (see Supplementary Material Section  I.7.3). Consistency with the experimental data was achieved by assuming that four out of the 68 fiber modes were leaky modes, with a transmission of $t_{0}=0.75$, corresponding to an average loss of less than 0.015 per channel (Fig. 2, purple curve). This assumption aligned with a direct measurement of the transmission matrix of a similar fiber without reflective coatings (see Supplementary Material, Section I.5). While neither model perfectly matches the experimental PDF, both indicate that the degree of control over the MMFC channels is high and that the overall loss is low.
@@
The MMFC platform, along with the high degree of control it provides over the incident and detected modes, enables us to selectively excite transmission eigenchannels and study their spatial structure. To excite an open channel of the system, we tailored the incident wavefront to match the first transmission eigenchannel of the transmission matrix, namely the input channel $\vec{v}_{1}$ with the highest transmission rate $\tau_{1}$. Since mode mixing in the MMFC is polarization-dependent (see Fig. 1b), the incident field corresponding to $\vec{v}_{1}$ consisted of two different polarization components. However, as the SLM can shape only one polarization at a time, we first sent the horizontal polarization component $\vec{v}_{1}$ into the fiber and measured the output field $\vec{\psi}_{1,H}$. At the output, both polarization components were measured simultaneously, so $\vec{\psi}_{1,H}$ includes both components, with the subscript $H$ indicating the input polarization. Next, we rotated the input polarization state by $90^{\circ}$, sent the vertical component of $\vec{v}_{1}$, and measured the output field $\vec{\psi}_{1,V}$. Finally, we coherently combined the measured output fields to obtain the total output field for simultaneous excitation of both input polarizations: $\vec{\psi}_{1}=\vec{\psi}_{1,H}+\vec{\psi}_{1,V}$, as depicted in Fig. 5.
@@
Figure 5 shows that the spatial distribution of the output fields $\vec{\psi}_{1,H}$ and $\vec{\psi}_{1,V}$ are highly correlated, despite their orthogonal input polarizations. This spatial overlap enables efficient constructive interference, which is required for open channels with near-unity transmission rates. Similarly, multiple round trips in the cavity are expected to interfere constructively at the fiber output. In a perfect fiber without mode mixing, guided modes maintain their transverse shape and open channels correspond to the guided modes whose propagation constants satisfy constructive interference at the output of the fiber. However, in the presence of mode mixing, none of the modes retain their transverse shape preventing good spatial overlap of multiple round trips. Indeed, when we coupled the guided modes of the fiber (LPlm) to the MMFC, the highest transmission rate we observed was $\tilde{\tau}_{LP}=0.17$, far below $\tilde{\tau}_{1}$. A unique feature of MMFC open channels is that they preserve their transverse shape after each round trip despite mode mixing (see Supplementary Material Section I.7.1). While we cannot directly probe the transverse shape after each round trip, we observed indirect evidence of this feature by tuning the wavelength of the laser by a few femtometers. For open channels, the intensity quickly dropped, while the transverse shape stayed constant. This suggests that the wavelength detuning was too small to change the interference between the modes, resulting in similar output intensity patterns, but large enough to change the relative phase between different round trips.
@@
The high degree of control was enabled by the unique properties of MMFCs, where the channels remain confined to the fiber core with a finite angular spread set by the numerical aperture, making them easily addressable with an SLM. This control enabled us to investigate the spatial structure of transmission eigenchannels, revealing that for channels with near-unit transmission rates, orthogonal input polarization components are highly correlated. This correlation underscores their origin in optimal constructive interference and long dwell times.

### v2

@@ Abstract
The transport of light in disordered media is governed by open transmission channels, which enable nearly complete transmission of the incident power, despite low average transmission. Extensively studied in diffusive media and chaotic cavities, open channels exhibit unique properties such as universal spatial structure and extended dwell times. However, their experimental study is challenging due to the large number of modes required for control and measurement. We propose a multimode fiber cavity (MMFC) as a platform to explore open channels. Leveraging mode confinement and finite angular spread, MMFCs enabled full channel control, yielding an 18-fold power enhancement in experiment by selectively exciting an open channel with a transmission rate of $0.90\pm 0.04$. By analyzing 100 transmission matrices of MMFC realizations, we observed a bimodal transmission eigenvalue distribution, indicating high channel control and low losses. The scalability of MMFCs, combined with long dwell times and potential for nonlinear phenomena, offers new opportunities for studying complex wave transport.
@@
Transmission through multimode optical fibers shares similarities with scattering samples, as both exhibit random mode coupling [18, 19, 20, 21]. A key advantage of multimode fibers is that all their modes can be controlled using an SLM [22, 23]. In standard optical fibers, open channels have little significance since backscattering is negligible. However, by introducing reflective coatings to create a multimode fiber cavity (MMFC), most incident light is backreflected and the concept of open channels becomes applicable. Without mode mixing, an open channel of an MMFC simply corresponds to a guided mode whose propagation constant matches the Fabry–Pérot resonance condition. However, in the presence of strong mode mixing, the existence and properties of open channels ̵ֿare more intricate.
@@
Open channels are found by measuring the transmission matrix of the sample $T$, which relates the input and output fields by $\vec{\psi}_{out}=T\vec{\psi}_{in}$ [26]. The singular value decomposition of $T$, or equivalently, the eigenvectors of the Hermitian matrix $T^{\dagger}T$, define the transmission eigenchannels of the sample $T^{\dagger}T\vec{v}_{n}=\tau_{n}\vec{v}_{n}$, where the eigenvalues $\tau_{n}$ correspond to the transmission rates. To selectively excite the most open channel, the wavefront of the input field is tailored to match the wavefront corresponding to the eigenchannel $\vec{v}_{1}$ with the highest transmission rate $\tau_{1}$. The two key ingredients for realizing open channels are, therefore, coherent detection of the light transmitted through the fiber and precise control over the amplitude and phase of the incident wavefront.
@@
To experimentally study the transmission eigenchannels of an MMFC, we used a one-meter long step-index fiber with a core diameter of $25\,{\mathrm{\mu m}}$ and a numerical aperture of NA = 0.1, coated with reflective coatings with reflectivity $\rho=0.88$. We placed the MMFC in one arm of a Mach-Zehnder interferometer (Fig. 1a and Fig. S1). A tunable laser ($\lambda=632\,\mathrm{nm}$) with a long coherence length ($>$ 100 m) enabled interference of multiple round trips in the cavity. The amplitude and phase of the incident field were tailored using a phase-only spatial light modulator (SLM) in combination with a spatial filter for amplitude modulation [27] (see Supplementary Material, Section I.1). The transmitted field was interfered with a reference beam in an off-axis holography configuration, and both output polarizations were imaged onto a camera. We normalized the reconstructed output field so that its total power matched the output power measured by a calibrated photodiode monitoring the transmitted light (see Supplementary Material Fig. S1 and Section I.2).
@@
The eigenvalues obtained by diagonalizing $T^{\dagger}T$ are presented in Fig. 2. We observed five channels with transmission rates higher than four times the average transmission, $\langle\tau\rangle=0.064\pm 0.001$. The presence of multiple channels with transmission rates exceeding $4\langle\tau\rangle$ demonstrates significant control over the incident wavefront. In contrast, partial control results in a transmission rate distribution that rapidly converges to the Marčenko-Pastur distribution, for which the maximal transmission rate for square matrices is $4\langle\tau\rangle$ [26, 28]. Additionally, we observed a consistent decrease in the eigenvalues up to a pronounced gap at the 67th eigenvalue. This suggests that the MMFC supports 67 modes, which is largely in line with the measured transmission matrix of the fiber without dielectric coatings, exhibiting 34 guided modes and two leaky modes per polarization (see Supplementary Material Section I.5).
Figure 2: Transmission rates $\tau_{n}$ obtained by diagonalizing $T^{\dagger}T$ for an experimentally measured transmission matrix $T$. Only the 70 most significant transmission rates are presented. The gap between channels 67 and 68 indicates that the MMFC supports 67 guided modes. The dashed horizontal line marks the average transmission $\langle\tau\rangle=0.064\pm 0.001$. The uncertainties are in the order of $1\%$, smaller than the marker size (see Supplementary Material Section I.4)
@@
Figure 4: Transmission matrix and mode mixing in a multimode fiber cavity (MMFC). (a) Schematic of the MMFC: Each round trip is described by $\hat{r}_{1}T_{0}^{\mathsf{T}}\hat{r}_{2}T_{0}$, where $T_{0}$ and $T_{0}^{\mathsf{T}}$ are the transmission matrix of an uncoated fiber and its transpose, respectively, and $\hat{r}_{1},\hat{r}_{2}$ denote the matrices describing the reflection from the facets. (b) Typical measured transmission matrix $T_{0}$ of a one-meter long uncoated stepped-index fiber supporting $N=34$ modes per polarization. The matrix exhibits weak mode coupling, with 0.8 of the energy concentrated in the diagonal blocks of the degenerate modes. (c) Numerically computed reflection matrix $\hat{r}_{1}$, assuming a facet angle of $\theta=5\times 10^{-3}$ rad, obtained by decomposing a linear phase tilt into the fiber modes. (d) Computed transmission matrix of the MMFC using Eq. (1), with for $T_{0}$ from (b) and $r_{i}$ from (c).
@@
The slight deviation measured and model PDFs can be used to estimate the MMFC channel control, as incomplete wavefront control or losses typically reduce the number of open channels. We therefore considered two simplified single-parameter models. In the incomplete channel control (ICC) model, we assumed a lossless MMFC, while allowing for an incomplete measurement of its transmission matrix (see Supplementary Material Section I.7.2 and Ref.[28]). The best agreement between the ICC model and the experimental data was obtained by assuming full control over all incident channels and 95% control over the output channels (Fig. 3, orange curve). In the loss model, we assumed the MMFC is lossy due to leaky fiber modes but that its transmission matrix was perfectly measured (see Supplementary Material Section  I.7.3). Consistency with the experimental data was achieved by assuming that the four highest-order modes of the fiber were leaky modes with a transmission of $t_{0}=0.75$, corresponding to an average loss of less than 0.015 per channel (Fig. 3, purple curve). This assumption aligned with a direct measurement of the transmission matrix of a similar fiber without reflective coatings (see Supplementary Material, Section I.5). Neither model perfectly fits the data, but both suggest high channel control and low overall loss.
@@
The MMFC platform’s high control enables selective excitation of transmission eigenchannels and study of their spatial structure. To excite an open channel of the system, we tailored the incident wavefront to match the first transmission eigenchannel of the transmission matrix, namely the input channel $\vec{v}_{1}$ with the highest transmission rate $\tau_{1}$. Since mode mixing in the MMFC is polarization-dependent (see Fig. 1b), the incident field corresponding to $\vec{v}_{1}$ consisted of two different polarization components. However, as the SLM can shape only one polarization at a time, we first sent the horizontal polarization component $\vec{v}_{1}$ into the fiber and measured the output field $\vec{\psi}_{1,H}$. At the output, both polarization components were measured simultaneously, so $\vec{\psi}_{1,H}$ includes both components, with the subscript $H$ indicating the input polarization. Next, we rotated the input polarization state by $90^{\circ}$, sent the vertical component of $\vec{v}_{1}$, and measured the output field $\vec{\psi}_{1,V}$. Finally, we coherently combined the measured output fields to obtain the total output field for simultaneous excitation of both input polarizations: $\vec{\psi}_{1}=\vec{\psi}_{1,H}+\vec{\psi}_{1,V}$, as depicted in Fig. 5.
@@
Figure 5 shows that the spatial distribution of the output fields $\vec{\psi}_{1,H}$ and $\vec{\psi}_{1,V}$ are highly correlated, despite their orthogonal input polarizations. This spatial overlap enables efficient constructive interference, which is required for open channels with near-unity transmission rates. Similarly, multiple round trips in the cavity are expected to interfere constructively at the fiber output. In a perfect fiber without mode mixing, guided modes maintain their transverse shape and open channels correspond to the guided modes whose propagation constants satisfy constructive interference at the output of the fiber. However, in the presence of mode mixing, none of the modes retain their transverse shape preventing good spatial overlap of multiple round trips. Indeed, when we coupled the guided modes of the fiber (LPlm) to the MMFC, the highest transmission rate we observed was $\tilde{\tau}_{LP}=0.17$, far below $\tilde{\tau}_{1}$. MMFC open channels uniquely retain transverse shape across round trips despite mode mixing (see Supplementary Material Section I.7.1). While we cannot directly probe the transverse shape after each round trip, we observed indirect evidence of this feature by tuning the wavelength of the laser by a few femtometers. For open channels, the intensity quickly dropped, while the transverse shape stayed constant. This suggests that the wavelength detuning was too small to change the interference between the modes, resulting in similar output intensity patterns, but large enough to change the relative phase between different round trips.
@@
MMFCs enable high control by confining channels with a finite angular spread, making them easily addressable with an SLM. This control revealed strong correlations between orthogonal input polarizations in near-unit transmission channels, highlighting their origin in optimal constructive interference and long dwell times.